Adigirl29:

RSTU is a parallelogram. If m∠TUV = 78° and m∠TVU = 54°, explain how you can find the measure of ∠SRV. Show all steps of your work, and refer to any properties of triangles, parallelograms, or triangle congruency theorems as necessary to justify your response./Users/adamariposada/Desktop/02_10_prt2_g4_flvs.png

Adigirl29:

Adigirl29:

uh I couldn't see the picture can you send it again?

Florisalreadytaken:

i just confused myself for a second

1 attachment
Florisalreadytaken:

these 2 triangles are congruent, just that 1 of them is horizontally flipped

Adigirl29:

okay so what do I put for the answer exactly?

Florisalreadytaken:

first, we have to find \[V\hat T U\]

Florisalreadytaken:

because \[V\hat T U = S\hat R V\]

Adigirl29:

oh okay so u is 78 and v is 154 correct?

Florisalreadytaken:

54* yes

Adigirl29:

uh im lost lmao

Florisalreadytaken:

ok, so the formula we will use is \[V\hat T U= 180- (\color{chocolate}{V \hat U T} + \color{gold}{T\hat V U}) \]

Adigirl29:

mhm

Florisalreadytaken:

where \[ \color{chocolate}{V \hat U T} = 78\] and \[ \color{gold}{T\hat V U} = 58 \]

Florisalreadytaken:

thus, \[ V\hat T U= 180- (\color{chocolate}{78} + \color{gold}{54}) \]

Adigirl29:

ohhhhhh

Florisalreadytaken:

\[ V\hat T U= \color{lightskyblue}{48} \]

Florisalreadytaken:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Florisalreadytaken because \[V\hat T U = S\hat R V\] \(\color{#0cbb34}{\text{End of Quote}}\) and that is our answer :D

Shayla:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Florisalreadytaken thus, \[ V\hat T U= 180- (\color{chocolate}{78} + \color{gold}{54}) \] \(\color{#0cbb34}{\text{End of Quote}}\) 54?

Shayla:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Florisalreadytaken where \[ \color{chocolate}{V \hat U T} = 78\] and \[ \color{gold}{T\hat V U} = 58 \] \(\color{#0cbb34}{\text{End of Quote}}\) ??

Adigirl29:

Huh?

Adigirl29:

wasn't it 54 TVU??

Florisalreadytaken:

it is 54 -- mb

Shayla:

okay sorry

Florisalreadytaken:

but the rest is correct no worries!

Adigirl29:

ohhhh okay, thank you so so much! all I have is one question left ill post it soon, thank you!

Florisalreadytaken:

you got it mate!

Adigirl29:

wait isn't TUV 78?

Florisalreadytaken:

yes ?

Adigirl29:

then why VUT in the equation?

Florisalreadytaken:

because that is how congruent angles work lol

Adigirl29:

oh kk

Florisalreadytaken:

they are alternate exterior angles thats why : )

Adigirl29:

oh kk